OpenStudy (anonymous):

how would i solve this to tell if it is always, sometimes, or never true? 2/3x+4=2x

OpenStudy (lalaly):

\[\frac{2}{3x}+4=2x\] multiply both sides by 3x so equation becomes \[2+4(3x)=2x(3x)\]now solve for x

OpenStudy (anonymous):

it's sometimes. because if you put 3 into the equation, it works. but if you put 0 it doesn't.

OpenStudy (anonymous):

thank you so much! how would you solve these equations 1. (m-2)-5=8-2(m-4) and A=1/2h(b1+b2) solving for b2

OpenStudy (skullpatrol):

Use the idea of balance

OpenStudy (skullpatrol):

Whatever you do to one side you do to the other.

OpenStudy (skullpatrol):

@LostInSchool Are you still there?

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (skullpatrol):

You will always be @LostInSchool if you don't learn for yourself.

OpenStudy (anonymous):

so for quesiton 1. you move the numbers to one side. so it turns into (m-2)+2(m-4)=8+5 when you move the numbers to one side the signs in front of the numbers needs to be changed also. then it'll turn into 3m-10=13 3m=23